Solve for x: y=1/3(x+1)

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

y = (1/3)x + 1/3
(1/3)x = y - 1/3
x = (y - 1/3)/(1/3)
x = 3y - 1
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by stanbon(75887)

Solve for x: y=(1/3)(x+1)
-----
3y = x+1
x = 3y-1
--------------
Cheers,
Stan H.
-------------